OpenStudy (anonymous):

how do i solve for x when i have cos^3x=2?

OpenStudy (phi):

does that say \[ \cos^3(x) = 2 \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

you could re-write the equation by taking the cube root of both sides to get \[ \cos(x) = \sqrt[3]{2} \] what is the cube root of 2 ?

OpenStudy (phi):

do you have a calculator ?

OpenStudy (anonymous):

1.2599

OpenStudy (phi):

ok, that looks good. Do you know what the biggest the cos(x) can every be ?

OpenStudy (anonymous):

um i think the domain of cosx is -infinity to infinity

OpenStudy (phi):

yes, but what is its range ?

OpenStudy (anonymous):

but the graph of cosx is bounded by y= + and -1

OpenStudy (anonymous):

It would be faster if you put it in a calculator to find your answer. ^_^

OpenStudy (phi):

cosx is bounded by y= + and -1 yes, cos(x) ≤ 1 which means it can never by 1.2599

OpenStudy (anonymous):

so the equation is false?

OpenStudy (phi):

there are no real solutions to this equation.

OpenStudy (anonymous):

i'm confused with the x in cosx and "bounded" referring to the y values

OpenStudy (phi):

if you plot cos(x) you start with y = cos(x) then for every x, you figure out what y is, and plot the point (x,y) you will get a cosine curve. it looks like this

OpenStudy (phi):

where did you get your equation \[ \cos^3(x) = 2 \]?

OpenStudy (anonymous):

i was doing a second derivative and ended up with that

OpenStudy (phi):

what was the original problem? Maybe you took a wrong turn to get that result.

OpenStudy (anonymous):

f=sinx=tanx i double checked my work and it seems right

OpenStudy (anonymous):

cos^3 x = 2 was just part of the f"

OpenStudy (anonymous):

i mean f=sinx-tanx

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

yes, OOOPS

OpenStudy (anonymous):

but there was no real answer so yeah

OpenStudy (anonymous):

thank phi

OpenStudy (anonymous):

why do you have to take second derivative to solve for x?

OpenStudy (anonymous):

and it's a function, how to "solve for x"

OpenStudy (anonymous):

oh, i didn't mean that. i'm testing for concavity so that's why i took the second derivative of f and then solved for x

OpenStudy (phi):

so that is f(x) = sin(x) - tan(x) and you want to solve for when f(x) = 0 ? or you want to plot f(x) ?

OpenStudy (anonymous):

my main goal is to plot f(x)

OpenStudy (phi):

it has zeros when sin(x) - sin(x)/cos(x) = 0 --> sinx cosx - sinx = 0 --> sinx(cosx - 1) = 0 --> sinx= 0 or cosx = 1 it goes to infinity (sign depends on sinx) when cos(x)=0

OpenStudy (phi):

the first derivative f'(x) = cos(x) - 1/cos^2(x) = 0 or cos^3(x) - 1 = 0 --> cos^3(x) = 1 which probably means you have an inflection point at x=0

OpenStudy (phi):

the 2nd derivative is zero when -sin(x) - 2 sin(x) cos^-3(x) =0 sin(x) ( cos^3(x) +2) = 0 that is zero when sin(x)=0 at x=0 so it looks like x=0 is an inflection point.